OpenStudy (anonymous):

Which of the following equations could be used to solve for the tenth term of the following sequence? 15, 13, 11, 9, ... A(10) = 15 + 10(-2) A(10) = 15 + 9(-2) A(10) = 15 + 9(2) A(10) = 15 + 10(2) how do

OpenStudy (anonymous):

The formula is A(n) = a1 + (a2 - a1)(n - 1) where a1 is the first term and a2 is the second term.

OpenStudy (anonymous):

thats what im getting. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

str99, now you plug in n = 10 and evaluate

OpenStudy (anonymous):

oh so its A?

OpenStudy (anonymous):

A or D right?

OpenStudy (anonymous):

A(10) = 15 + (13 - 15)(10 - 1) = 15 + 9(-2)

OpenStudy (anonymous):